OpenStudy (anonymous):

x^2+2x+1=0 solution set?

OpenStudy (anonymous):

-1 and -1

OpenStudy (anonymous):

two roots

OpenStudy (er.mohd.amir):

this (x+1)^2 so ans x=-1,-1

OpenStudy (anonymous):

What about x^2+4x+4=0?

OpenStudy (anonymous):

x^2 +2x + 1 =0 (x+1)^2 = 0 so x = -1 and -1

OpenStudy (er.mohd.amir):

(x+2)^2=0 x=-2,-2

OpenStudy (anonymous):

x^2 + 4x +4 = 0 (x+2)^2 = 0 so x = -2 and -2

OpenStudy (anonymous):

x^2-14x+49?

OpenStudy (anonymous):

(x - 7) ^2 = 0 so x = 7 and 7 are the required roots

OpenStudy (anonymous):

x^2+16x+64 Would be 8 right?

OpenStudy (anonymous):

(x+8)^2 = 0 so x = -8

OpenStudy (anonymous):

note -8 not 8

OpenStudy (anonymous):

x^2+24x+144=0 is -12?